OpenStudy (anonymous):

2. Consider the differential equation dy dt = t y . (a) Show by substitution into the equation that y = t is a solution, satisfying the initial condition y(1) = 1. Over what range of t is this solution valid? (b) Find the solution satisfying the intitial condition y(0) = 2.

OpenStudy (anonymous):

im not sure what the equation is is it dy/dt = ty ?

OpenStudy (anonymous):

t-y

OpenStudy (anonymous):

sorry t/y

OpenStudy (ash2326):

dy/dt =t/y ydy= tdt integrate both sides y^2/2=t^2/2+c y^2=t^2+C given y(1)=1 1=1+c so c=0 y^2=t^2 y=t this is valid for t>=0 2nd part y(0)=2 2=0+c so y^2=t^2+2